OpenStudy (anonymous):

To get a chance to win a car, you must guess then number of keys in a jar to within 5 of the actual number. Those who are within this range are given a key to try in the ignition of the car. Suppose there are 587 keys in the jar. Write and solve and equation to determine the highest and lowest guesses that will give contestants a chance to win the car.

myininaya (myininaya):

so we need to be with in 5 units of 587 so we can have 587+5 and 587-5 or anything in between so we need x to be btw 587+5 and 587-5 like so 587-5<x<587+5

myininaya (myininaya):

582<x<592

myininaya (myininaya):

but it says to write an eqaution that will give the highest and lowest guesses

OpenStudy (anonymous):

Thank you so much! I thought that was the answer but i wasnt sure because it said equation and not inequality.

OpenStudy (anonymous):

Yeah that is why i didnt think that was right. Maybe i will jst have to write an interval notation with it.

myininaya (myininaya):

let me think

myininaya (myininaya):

---|----|-----|----- 582 587 592 --|----|----|---- <- we can make an equation for this one |x|=5 right? -5 0 5 so lets think about the fist one we are not centered at 0 we are centered at 587 so we have |x-587|=5

myininaya (myininaya):

the equation to obtain the highest and lowest is |x-587|=5

myininaya (myininaya):

Check: |x-587|=5 x-587=5 or x-587=-5 x=587+5 x=-5+587 x=592 x=582

myininaya (myininaya):

:)

myininaya (myininaya):

do you have any question?

OpenStudy (anonymous):

No. That helped me completely. Thank You so much!